translate to a system of equations and solve.

roberta is 25 years older than her daughter, Cindy. In four years, Roberta will be twice as old as Cindy. how old are they now?

Answer by xcentaur(357)

let the age of roberta be x
let the age of cindy be y


x=25+y........[1] (given)


in four years,
robertas age=(x+4)
cindys age=(y+4)
given that roberta will be twice as old as cindy
therefore,
x+4=2(y+4)........[2]


x=25+y...[1]
x+4=2(y+4)
x+4=2y+8
x=2y+8-4
x=2y+4........[3]


from eqns 1 and 3 we get
x=25+y=2y+4
25+y=2y+4
25-4=2y-y
y=21
x=25+21=46


Hence their current ages are:
Roberta:46
Cindy:21